# Learning to detect continuous gravitational waves: an open data-analysis competition

### v1

@@ Abstract
We report results of a public data-analysis challenge, hosted on the open data-science platform Kaggle, to detect simulated continuous gravitational-wave signals. These are weak signals from rapidly spinning neutron stars that remain undetected despite extensive searches. The competition dataset consisted of a population of CW signals using both simulated and real LIGO detector data matching the conditions of actual CW searches. The competition attracted more than 1,000 participants to develop realistic CW search algorithms. We describe the top 10 approaches and discuss their applicability as a pre-processing step compared to standard CW-search approaches. For the competition’s dataset, we find that top approaches can reduce the computing cost by 1 to 3 orders of magnitude at a $2\%$ dismissal probability. Additionally, the competition drove the development of new GPU-accelerated detection pipelines and extended CW-inspired statistics to other domains. We release the associated dataset, which constitutes the first open standardized benchmark for CW detection, to enable reproducible method comparisons and to encourage further developments toward the first detection of these elusive signals.
@@ 1 Introduction
Continuous gravitational waves (CWs) [1, 2, 3, 4] are long-duration gravitational-wave (GW) signals expected from rapidly-spinning non-axisymmetric neutron stars (NSs). Although currently undetected, their detection would, shed light on the extreme physics of NSs, allow us to understand their population in our galaxy, or even establish a new probe to direct and indirect detection of dark matter.
CW searches can be classified according to their assumptions on the expected source [5]. Targeted searches, for example, assume the CW emission of a pulsar fully matches its electromagnetic emission; this allows for very sensitive searches at the cost of not being robust to slight mismatches of the signal. On the other end of the spectrum, blind searches attempt to detect a CW signal from unknown sources; due to their parameter-space breadth, these kinds of searches cannot make use of matched-filtering and instead resolve to semicoherent methods to trade sensitivity for computing cost [6]. Incidentally, these approaches make blind searches more robust to un-modelled physics such as NS glitches [7] or spin-wandering [8, 9].
@@
Machine learning applications [35] have demonstrated significant accelerations in search [36, 37, 38, 39] and parameter-estimation [40, 41, 42, 43] workflows targeting compact binary coalescences compared to standard approaches [44, 45, 46]. Due to the comparatively broader parameter space, applications to blind CW signals have reported limited success [47, 48, 49, 50, 51], with most successful applications focusing on accelerating specific computational steps [52] or increasing the robustness of a search to unmodelled physics [53, 54, 55].
In this work, we present the results of an open data-analysis competition to spark new approaches and algorithmic developments in CW searches [56]. This competition delivered a comparison of methods beyond what within-domain mock data challenges are able to prospect [57, 58, 59, 60] and allowed us to understand the complications and computational advantages achievable by off-the-shelf approaches from beyond the community. We release the dataset of the competition [61] to provide the first publicly-available benchmark targeted towards the detection of CW searches.
We also summarize the results of the first Kaggle competition aimed at detecting GW signals from compact binary coalescences [62] in A.
@@ 2 Leveraging open competitions for gravitational-wave research
Open data-analysis competitions have demonstrated significant value in advancing scientific research across multiple domains, such as medical imaging [63, 64, 65], time-series forecasting [66], or particle physics [67], often producing solutions that outperform existing standards. These successes stem from competitions’ ability to attract diverse problem-solving approaches, encourage rapid iteration, and provide standardized benchmarks for method comparison [68].
The gravitational-wave community has already embraced crowd-sourcing approaches through platforms, such as Zooniverse [69], to understand transient noise in interferometric detectors [70, 71, 72], or Kaggle [73] to search for gravitational-wave signals from compact binary coalescences [62]. This approach can be understood as an evolution of the well-established mock data challenges conducted by the GW community [57, 58, 59, 60], where the aim has shifted from comparing existing pipelines to exploring new strategies altogether. The overall experience in GW citizen-science so far suggests non-experts, when given a proper training framework, can deliver solutions beneficial to the GW community [74, 39, 75].
The problem of detecting CW signals presents an ideal case for an open competition: First, the problem can be framed as a well-defined classification task [76, 77, 78], which provides a simple criterion to rank different strategies against each other. Second, synthetic data can be easily generated, either from first principles or through well-established software packages [79, 80, 81]. Third, since large-scale searches can be understood as the aggregation of many small-scale searches [29], competitors can tackle a simulation of the real problem using standard hardware. Finally, despite the existing experience in the field, large portions of the search-pipeline configuration space where novel improvements may be found remain yet to be explored.
Building on previous positive results [62], we released our competition on Kaggle [73], a well-established data-analysis platform powered by Google with proven track record in scientific applications and more than a million active users spread across vast domains of data science and engineering.
@@ 3 Dataset description
We seek a data-analysis strategy compatible with a blind CW search; that is, a pipeline which classifies narrow frequency bands of LIGO/Virgo/KAGRA [82, 83, 84, 85] data according to whether they contain a signal or not. We take the deep searches conducted using GCT [86] and Weave [87] as a gold standard for our competition. These pipelines use highly-efficient implementations of the semicoherent $\mathcal{F}$-statistic [20, 21, 23, 24, 22] to deliver the most sensitive all-sky search results across their explored parameter space. Their sensitivity is effectively bounded by the available computing resources, which amount to about 100 million CPU hours or 1 million GPU hours, depending on the specific implementation. By framing CW detection as a binary classification problem, we aim to discover whether alternative approaches could compete with or complement these well-established strategies.
@@ 3.1 Data format
We constructed a dataset using Short Fourier Transforms [88] (SFTs), a lossless data format used by actual blind CW searches [26, 23, 24, 53, 28, 25, 29]. This choice provides the participants with the same information available to CW data analysts and simplifies the transition of a successful solution into a production stage.
Given a set of $N$ time-domain data samples $x_{\alpha}[j],j=0,\dots,N-1$ starting at an epoch $t_{\alpha}$ sampled with a period $\Delta t$, we define an SFT following [88]:
@@
The competition dataset consisted of 8,000 test samples divided into signal and noise samples in equal parts. Additionally, 600 training samples were provided to the competitors as a quick entrypoint to the competition. The dataset was made available on the Kaggle platform, and adds up to $\sim\!0.5\,\mathrm{TB}$. Additionally, we supplied the tools to generate simulated CW data through the PyFstat package [81], which provides a Python interface to the LIGO/Virgo/KAGRA Algorithm Library [79, 80].
The dataset contains both noise-only samples and software-simulated CW signals, using both simulated Gaussian noise and data slices from the third observing run of the Advanced LIGO detectors [89, 90, 91]. The chosen frequency band ($0.2\text{\,}\mathrm{Hz}$ or $0.2\text{\,}\mathrm{Hz}$\times T_{\textrm{SFT}}=360$ frequency bins) is compatible with those analyzed by current blind searches. The sample duration, on the other hand, falls on the shorter end, as most searches analyse at least 6 months of data. This choice is due to the limited storage available on the Kaggle platform. Regardless, the duration is long enough to contain the main distinctive features of CW signals, namely a Doppler shift on the time-scale of months combined with a daily amplitude modulation.
@@ 3.2 Noise distribution
Each data sample in the competition’s dataset consists of a two sets of Short Fourier Transforms, representing the H1 (Hanford) and L1 (Livingston) LIGO detectors, spanning a total time of 4 months and a frequency band of $0.2\text{\,}\mathrm{Hz}$ with $T_{\textrm{SFT}}=$1800\text{\,}\mathrm{s}$\$. To better simulate the conditions of a real search, we introduce gaps in the data. As a result, SFTs are not necessarily contiguous (i. e. $t_{\alpha+1}-t_{\alpha}\geq T_{\textrm{SFT}}$).
@@
To generate a data sample, we fix the total duration at $T_{\textrm{span}}=4\;\mathrm{month}$ and sample gap/science-mode durations alternatively. The first segment is sampled from the gap distribution and added to the starting time of O3 to avoid fixing the starting time. This process is repeated for both interferometers independently. By construction, the resulting data samples have a “duty cycle” (fraction of science-mode data) of
which is comparable with realistic observing runs [93, 94].
To generate a data sample containing real noise, we sample a subset of the full observing run amounting to $T_{\textrm{span}}=4\;\mathrm{month}$ following the distributions in Fig. 1. To avoid cross-correlation of strong instrumental artifacts with actual Advance LIGO data, we shift the frequency labels a random amount within $\pm 4\text{\,}\mathrm{Hz}$ and add random Gaussian noise with 0.1% of the data’s ASD.
@@ 3.3 Signal distribution
We simulate a source population compatible with that assumed in blind searches, as summarized in Table 1. To generate a signal sample, we draw a set of parameters $(\mathcal{A},\lambda)$ from the population, we simulate noise as described in Sec. 3.2 for both the H1 and L1 LIGO detectors, and use MakeFakeData_v5 [79] to add a simulated signal into the noise. This software simulates a CW signal as observed by a ground-based interferometric detector, adds Gaussian noise with a given ASD, and returns a the corresponding SFTs with the specified $T_{\textrm{SFT}}$.
The sky position $\hat{n}$ is drawn uniformly across the celestial sphere; the distribution over the initial phase and orientation angles $(\cos\iota,\psi)$ follows that of an isotropically oriented-population. The spindown $f_{1}$ is drawn from a log-uniform distribution along $10^{-8},-10^{-12}]$ with 90% probability and $10^{-12},10^{-9}]$ with a 10% probability. This split accounts for possible spin-ups due to accretion, proper motion, or the evaporation of a boson cloud [20, 8, 95]
The frequency $f_{0}$ is sampled uniformly along $50,500]\,\mathrm{Hz}$. The maximum value of $500\text{\,}\mathrm{Hz}$ is chosen so that the CW signal does not drift more than one frequency bin within an SFT. The bandwidth of a CW signal depends on its frequency and sky location, as it is largely dominated by the Doppler effect due to the orbital motion of the detector around the Sun. For the parameter space considered in this challenge, the maximum bandwidth is about $0.05\text{\,}\mathrm{Hz}$. To complete a band of $0.2\text{\,}\mathrm{Hz}$, we attach noise data to both sides of the signal’s spectrum so it is contained completely within the data sample. Also, to avoid biasing the dataset, the proportion of data attached to each side of the signal’s frequency band is chosen randomly so that the signal is not centred.
@@
Blind CW searches usually quote their results in terms of population-based upper limits at a certain dismissal probability (usually 5% to 10%) on a constant-$h_{0}$ population [3, 5]. The false-alarm probability of a search is usually not computed. This complicates the comparison of Kaggle solutions to realistic searches on equal footing. Instead, we will base our comparisons on computing cost; concretely, we will assess the capability of Kaggle solution to discard broad parameter-space regions in such a way that signals detectable by a blind CW search are not discarded. To do so, it will suffice to re-scale the sensitivity estimates reported by the two reference searches to the dataset duration used in the competition.
The two reference searches we use in this challenge, based on the GCT [86] and Weave [87] pipelines, report their sensitivity estimates in terms of the average amplitude $\langle h_{0}\rangle$ at which 90% of an isotropically-oriented uniformly-sky-distributed signals would be detected. We relate this quantity to the corresponding averaged SNR as [96]
@@
These pipelines operate by computing the semicoherent $\mathcal{F}$-statistic [20, 21]. Broadly, the dataset is split into segments with a duration $T_{\textrm{coh}}$ which are analysed independently using the $\mathcal{F}$-statistic, and then the results are combined to compute the semicoherent $\mathcal{F}$-statistic [23, 24, 22].
@@
With this information, we choose to describe the SNR distribution using a Gamma distribution with shape parameter $k=5$ and scale parameter $\theta=9$, shown in Fig. 2. This distribution is such that about 20% of the signals are above the sensitivity of GCT and about 50% of the signals are above the sensitivity of Weave. About 6% of the signals display visible features ($\rho>80$).
@@
GCT [86] 60 $1.3\times 10^{-25}$ 47 61
Weave [87] 240 $1.0\times 10^{-25}$ 36 47
@@
Figure 2: Empirical cumulative distribution of $\rho\sim\Gamma(5,9)$. Vertical lines denote the re-scaled sensitivity of different search pipelines as discussed in the main text.

### v2

@@ Abstract
We report results of a public data-analysis challenge, hosted on the open data-science platform Kaggle, to detect simulated continuous gravitational-wave signals (CWs). These are weak signals from rapidly spinning neutron stars that remain undetected despite extensive searches. The competition dataset consisted of a population of CW signals using both simulated and real LIGO detector data matching the conditions of actual CW searches. The competition attracted more than 1,000 participants to develop realistic CW search algorithms. We describe the top 10 approaches and discuss their applicability as a pre-processing step compared to standard CW-search approaches. For the competition’s dataset, we find that top approaches can reduce the computing cost by 1 to 3 orders of magnitude at a false-dismissal probability comparable to standard CW searches. Additionally, the competition drove the development of new GPU-accelerated detection pipelines, which facilitated their adoption in other areas of gravitational-wave data analysis. We release the associated dataset, which constitutes the first open standardized benchmark for CW detection, to enable reproducible method comparisons and to encourage further developments toward the first detection of these elusive signals.
@@ 1 Introduction
Continuous gravitational waves (CWs) are long-duration gravitational-wave (GW) signals expected from rapidly-spinning non-axisymmetric neutron stars (NSs) [Riles:2022wwz]. Although currently undetected,their detection would shed light on the extreme physics of NSs [Haskell:2023yrv], allow us to understand their population in our galaxy [Pagliaro:2023bvi, Hua:2023aff], or even establish a new probe to direct and indirect detection of dark matter [Miller:2025yyx].
CW searches can be classified according to their assumptions on the expected source [Wette:2023dom]. Targeted searches, for example, assume the CW emission of a pulsar is phase-locked to its electromagnetic emission; this allows for very sensitive searches using matched filtering. Blind searches, on the other hand, attempt to detect a CW signal from unknown sources; due to the sheer size of their parameter-space, these kinds of searches cannot make use of matched filtering.
A standard CW search operates by filtering data against a set of signal models, usually referred to as templates, which are parameterised by their amplitude and phase-evolution parameters (see Sec. 3.3). The result of filtering data against a specific signal model is a detection statistic, which is then used to establish the significance of a candidate. The specific detection statistic in use depends on search assumptions and available computing budget, and has a direct impact on sensitivity [Tenorio:2021wmz].
Fully-coherent searches typically use the $\mathcal{F}$-statistic [Jaranowski:1998qm, Cutler:2005hc], which tracks the phase of a CW signal accounting for possible amplitude and frequency modulations due to the orbital motion of GW detectors. This approach is unfeasible for broad parameter spaces as the number of required signal templates quickly becomes computationally prohibitive [Wette:2015lfa]. Semicoherent detection statistics, on the other hand, follow the frequency evolution of a signal and only track its phase evolution along short segments. Effectively, a semiocherent statistic accumulates coherent statistics (such as segment-wise $\mathcal{F}$-statistics, power spectra [soap, Tenorio:2024jgc], or the binarized power [Krishnan:2004sv, 2014PhRvD..90d2002A]) following the signal’s frequency evolution. These statistics impose milder constraints and require less templates to cover a given parameter-space region, reducing the computing cost of a search, but also reduce the sensitivity as background noise is more likely to fit a signal template [Prix:2012yu]. Because of this, semicoherent searches are robust to unmodelled physics such as NS glitches [Ashton:2017wui] or spin-wandering [Mukherjee:2017qme, Carlin:2025sxm].
@@
Machine learning applications [Cuoco:2024cdk] have demonstrated significant accelerations in search [Gabbard:2017lja, Chua:2019wwt, Koloniari:2024kww, Nagarajan:2025hws] and parameter-estimation [Gabbard:2019rde, Dax:2021tsq, Williams:2021qyt, Williams:2023ppp] workflows targeting compact binary coalescences compared to standard approaches [Usman:2015kfa, Sachdev:2019vvd, Cornish:2020dwh]. Due to the comparatively broader parameter space, applications to blind CW signals have reported limited success [Dreissigacker:2019edy, Yamamoto:2020pus, Joshi:2025xdz, Cheung:2025lyf], with most successful applications focusing on accelerating specific computational steps [Modafferi:2023nzt] or increasing the robustness of a search to unmodelled physics [Bayley:2020zfa, Bayley:2022hkz].
In this work, we present the results of an open data-analysis competition to spark new approaches and algorithmic developments in CW searches [competition].
We constructed a dataset using the sensitivity of flagship searches such as GCT or Weave as a baseline reference to maximise the impact of winning solutions, and compared their performance to an agnostic CW search (SOAP) to gauge the impact of design choices made by different teams. This competition delivered a comparison of methods beyond what within-domain mock data challenges [Messenger:2015tga, Walsh:2016hyc, Schafer:2022dxv, Baghi:2022ucj] are able to prospect and allowed us to understand the complications and computational advantages achievable by off-the-shelf approaches from beyond the community. We release the dataset of the competition [dataset] to provide the first publicly-available benchmark targeted towards the detection of CW searches. Tools and tutorials to interact with this data are publicly available in Kaggle [kaggle_post] and the PyFstat repository [pyfstat_github].
The paper is structured as follows: In Sec. 2 we introduce the rationale behind this CW-search data-analysis competition. In Sec. 3 we introduce the competition’s dataset. In Sec. 4 we discuss the top-scoring solutions of the competition. In Sec. 5 we interpret these solutions in the context of standard CW searches. We finish by discussing these results in Sec. 6. We also summarize the results of the first Kaggle competition aimed at detecting GW signals from compact binary coalescences [first_challenge] in A.
@@ 2 Leveraging open competitions for gravitational-wave research
Open data-analysis competitions have demonstrated significant value in advancing scientific research across multiple domains, such as medical imaging [lungCompetition, BrainTumorPrediction, RNAprediction], time-series forecasting [forecasting], or particle physics [Calafiura:2018cdd], often producing solutions that outperform existing standards. These successes stem from competitions’ ability to attract diverse problem-solving approaches, encourage rapid iteration, and provide standardized benchmarks for method comparison [CompetitionsAreGold].
The gravitational-wave community has already embraced crowd-sourcing approaches through platforms, such as Zooniverse [zooniverse], to understand transient noise in interferometric detectors [Zevin:2016qwy, Zevin:2023rmt, Razzano:2022lgg], or Kaggle [kaggle] to search for gravitational-wave signals from compact binary coalescences [first_challenge]. This approach can be understood as an evolution of the well-established mock data challenges conducted by the GW community [Messenger:2015tga, Walsh:2016hyc, Schafer:2022dxv, Baghi:2022ucj], where the aim has shifted from comparing existing pipelines to exploring new strategies altogether. The overall experience in GW citizen-science so far suggests non-experts, when given a proper training framework, can deliver solutions beneficial to the GW community [Soni:2021cjy, Mackenzie:2025gut]. For example, the results of the first Kaggle competition [first_challenge] served as a basis to push forward our understanding of GW searches using machine learning [Nagarajan:2025hws].
The problem of detecting CW signals presents an ideal case for an open competition: First, the problem can be framed as a well-defined classification task [Prix:2009tq, Searle:2008jv], which provides a simple criterion to rank different strategies against each other. Second, synthetic data can be easily generated, either from first principles or through well-established software packages [lalsuite, swiglal, pyfstat]. Third, since large-scale searches can be understood as the aggregation of many small-scale searches [Tenorio:2024jgc], competitors can tackle a simulation of the real problem using standard hardware. Finally, despite the existing experience in the field, large portions of the search-pipeline configuration space where novel improvements may be found remain yet to be explored.
Building on previous positive results [first_challenge], we released our competition on Kaggle [kaggle], a well-established data-analysis platform powered by Google with proven track record in scientific applications and more than a million active users spread across vast domains of data science and engineering.
@@ 3 Dataset description
We seek a data-analysis strategy compatible with a blind CW search; that is, a pipeline which classifies narrow frequency bands of LIGO/Virgo/KAGRA [LIGOScientific:2014pky, VIRGO:2014yos, KAGRA:2018plz, KAGRA:2020tym] data according to whether they contain a signal or not. We take the sensitivity of deep searches conducted using GCT [eah_O2] and Weave [Wette:2021tbv] as a reference for our competition. These pipelines use highly-efficient implementations of the semicoherent $\mathcal{F}$-statistic [Pletsch:2009uu, Wette:2018bhc, Dunn:2022gai] to deliver the most sensitive all-sky search results across their explored parameter space. Their sensitivity is effectively bounded by the available computing resources, which amount to about 100 million CPU hours or 1 million GPU hours, depending on the specific implementation. By framing CW detection as a binary classification problem, we aim to discover whether alternative approaches could compete with or complement these well-established strategies. Note that the high computing cost of these searches complicates a direct comparison of methods; instead, we use the sensitivity of these pipelines to construct a signal distribution for the challenge, and after the competition discuss the advantages of the methods presented by the competitors.
@@ 3.1 Data format
We constructed a dataset using Short Fourier Transforms [SFTFormat] (SFTs), a data format used by actual blind CW searches. This choice provides the participants with the same information available to CW data analysts and simplifies the transition of a successful solution into a production stage.
Given a set of $N$ time-domain data samples $x_{\alpha}[j],j=0,\dots,N-1$ starting at an epoch $t_{\alpha}$ sampled with a period $\Delta t$, we define an SFT following [SFTFormat]:
@@
The competition dataset consisted of 8,000 test samples divided into signal and noise samples in equal parts. Additionally, 600 training samples were provided to the competitors as a quick entrypoint to the competition. The dataset was made available on the Kaggle platform, and adds up to $\sim\!0.5\,\mathrm{TB}$. Additionally, we supplied the tools to generate simulated CW data through the PyFstat package [pyfstat], which provides a Python interface to the LIGO/Virgo/KAGRA Algorithm Library [lalsuite, swiglal].
The dataset contains both noise-only samples and software-simulated CW signals, using both simulated Gaussian noise and data slices from the third observing run of the Advanced LIGO detectors [KAGRA:2023pio, O3aData, O3bData]. The chosen frequency band ($0.2\text{\,}\mathrm{Hz}$ or $0.2\text{\,}\mathrm{Hz}$\times T_{\textrm{SFT}}=360$ frequency bins) is compatible with those analyzed by current blind searches. The sample duration, on the other hand, falls on the shorter end, as most searches analyse at least 6 months of data. This choice is due to the limited storage available on the Kaggle platform. Regardless, the duration is long enough to contain the main distinctive features of CW signals, namely a Doppler shift on the time-scale of months combined with a daily amplitude modulation.
@@ 3.2 Noise distribution
All samples of the competition contain background noise. This was either simulated Gaussian noise or real Advanced LIGO noise from the third LVK observing run (O3) [KAGRA:2023pio, O3aData, O3bData].
Each data sample in the competition’s dataset consists of a two sets of Short Fourier Transforms, representing the H1 (Hanford) and L1 (Livingston) LIGO detectors, spanning a total time of $T_{\textrm{span}}=4\;\mathrm{month}$ and a frequency band of $0.2\text{\,}\mathrm{Hz}$ with $T_{\textrm{SFT}}=$1800\text{\,}\mathrm{s}$\$. To better simulate the conditions of a real search, we introduce gaps in the data. As a result, SFTs are not necessarily contiguous (i. e. $t_{\alpha+1}-t_{\alpha}\geq T_{\textrm{SFT}}$).
@@
For each interferometer, we draw gap/science-mode durations alternatively from the distributions in Fig. 1 until completing a span of $T_{\textrm{span}}$. The first segment is sampled from the gap distribution and added to the starting time of O3 to avoid fixing the starting time. As a result, the SFTs of different detectors have, in general, different starting timestamps. By construction, the resulting data samples for both detectors have a “duty cycle” (fraction of science-mode data) of
which is comparable with realistic observing runs [Buikema_2020, LIGO:2021ppb].
@@
To generate a data sample containing real noise, we instead sample SFTs from LVK O3 data. To prevent participants from identifying noise realisations using publicly available data, we shift the frequency labels by a random amount within $\pm 4\text{\,}\mathrm{Hz}$ and add Gaussian noise with 0.1% of the data’s ASD to mangle the numerical noise values but maintain real-noise features. This value was chosen empirically.
@@ 3.3 Signal distribution
We simulate a source population compatible with that assumed in blind searches, as summarized in Table 1. To generate a signal sample, we draw a set of parameters $(\mathcal{A},\lambda)$ from the population, we simulate noise as described in Sec. 3.2 for both the H1 and L1 LIGO detectors, and use MakeFakeData_v5 [lalsuite] to add a simulated signal into the noise. This software simulates a CW signal as observed by a ground-based interferometric detector, adds Gaussian noise with a given ASD, and returns a the corresponding SFTs with the specified $T_{\textrm{SFT}}$.
The sky position $\hat{n}$ is drawn uniformly across the celestial sphere; the distribution over the initial phase is uniform along $0,2\pi)$; orientation angles $(\cos\iota,\psi)$ are drawn following an isotropically-oriented population. The spindown $f_{1}$ is drawn from a log-uniform distribution along $10^{-8},-10^{-12}]\,\mathrm{Hz}/\mathrm{s}$ with 90% probability and $10^{-12},10^{-9}]\,\mathrm{Hz}/\mathrm{s}$ with a 10% probability. This split accounts for possible spin-ups due to accretion, proper motion, or the evaporation of a boson cloud [Jaranowski:1998qm, Mukherjee:2017qme, Zhu:2020tht]
The frequency $f_{0}$ is sampled uniformly along $50,500]\,\mathrm{Hz}$, which is a frequency range commonly explored by blind CW searches as it contains the most sensitive band of the Advanced LIGO detectors [Buikema_2020]. The bandwidth of a CW signal depends on its frequency and sky location, as it is largely dominated by the Doppler effect due to the orbital motion of the detector around the Sun. For the parameter space considered in this challenge, the maximum bandwidth is about $0.05\text{\,}\mathrm{Hz}$. To complete a band of $0.2\text{\,}\mathrm{Hz}$, we attach noise data to both sides of the signal’s spectrum so it is contained completely within the data sample. Also, to avoid biasing the dataset, the proportion of data attached to each side of the signal’s frequency band is chosen by drawing a uniform random number between 0 and 1 so that the signal is not systematically centred.
@@
Blind CW searches usually quote their results in terms of population-based upper limits at a certain false-dismissal probability (usually 5% to 10%) on a constant-$h_{0}$ population [Tenorio:2021wmz, Wette:2023dom]. The false-alarm probability of a search is usually not computed. This complicates the comparison of Kaggle solutions to realistic searches on equal footing. Instead, we will base our comparisons on computing cost; concretely, we will assess the capability of Kaggle solution to discard broad parameter-space regions in such a way that signals detectable by a blind CW search are not discarded. To do so, it will suffice to re-scale the sensitivity estimates reported by the two reference searches to the dataset duration used in the competition.
The two reference searches we use in this challenge, based on the GCT [eah_O2] and Weave [Wette:2021tbv] pipelines, report their sensitivity estimates in terms of the average amplitude $\langle h_{0}\rangle$ at which 90% of an isotropically-oriented uniformly-sky-distributed signals would be detected. We relate this quantity to the corresponding averaged SNR as [Wette:2011eu]
@@
These pipelines split the dataset into segments with a duration $T_{\textrm{coh}}$ which are analysed independently using the $\mathcal{F}$-statistic, and then the results are combined to compute the semicoherent $\mathcal{F}$-statistic.
@@
GCT [eah_O2] 60 $1.3\times 10^{-25}$ 47 61
Weave [Wette:2021tbv] 240 $1.0\times 10^{-25}$ 36 47
@@
Figure 2: Empirical cumulative distribution of SNR $\rho\sim\Gamma(5,9)$. Vertical lines denote the re-scaled sensitivity of different search pipelines as discussed in the main text.
With this information, we choose to describe the SNR distribution using a Gamma distribution with shape parameter $k=5$ and scale parameter $\theta=9$, shown in Fig. 2. This distribution is such that 20% of the signals are above the sensitivity of GCT and about 41% of the signals are above the sensitivity of Weave. 6% of the signals display visible features ($\rho>80$). We show an example CW signal at different SNR levels in Fig. 3.
Figure 3: Spectrograms of a sample containing 4 months of simulated Gaussian noise with a fiducial CW signal as observed by the H1 detector. The CW amplitude in the left (right) spectrogram is such that $\rho=350$ ($\rho=35$); the signal parameters are otherwise identical. As a result, the signal is visible in the left spectrogram but not in the right one. This data was generated without gaps for the sake of visualization.